OpenStudy (periodicninja):

can someone help me with explain what geometric and arithmetic sequences are and what the differences between them are?

OpenStudy (_malice_):

Well I believe one is geometric and the other is arithmetic.

OpenStudy (periodicninja):

@.Sam.

OpenStudy (periodicninja):

@Kainui

OpenStudy (jadeishere):

An arithmetic sequence is a sequence that creates a thing called common difference, whereas a geometric sequences creates the common ratio. A sequence is a set of numbers, arranged in a specific order. The two types of sequences rely on the consecutive numbers to remain in constant differences(arithmetic) , or constant ratios (geometric)

OpenStudy (agent0smith):

2, 4, 6, 8, who do we appreciate, is arithmetic

OpenStudy (jadeishere):

For example; 1, 2, 3, 4, 5, 6, 7, 8, 9, 10 2-1 = 1 3-2 = 1 4-3 = 1 5-4 = 1

OpenStudy (jadeishere):

Or, As he said

OpenStudy (periodicninja):

thanks!

OpenStudy (jadeishere):

No problem, can you determine the differences and similarities by yourself, or would you still like some help? :)

OpenStudy (periodicninja):

they are pretty much opposites, one is finding the differences of numbers and the other is finding similarities or constants. right?

OpenStudy (zootedyoungin):

is it because you got 3 medals?

OpenStudy (periodicninja):

shhhhh, dont be rude :(

OpenStudy (jadeishere):

Yes, while it's true one finds the difference where the other finds the ratio, there are some similarities! One could be that the difference and the ratio has to be between consecutive numbers, another is that they both have to remain constant in order to be considered a sequence.

OpenStudy (zootedyoungin):

I have trouble doing that

OpenStudy (kainui):

What's the area of this geometric figure  The total area is 2, cause each square has area 1, but if you look at that right square, I sliced it up into a half. Then I took the other half and cut it into a half again to get a fourth, then I continued on cutting... \[2=1+\frac{1}{2}+\frac{1}{4}+\frac{1}{8}+\frac{1}{16}+\cdots\] This is why they call it a geometric series, cause there's a geometric picture to go along with adding terms where each term is multiplied by a number, in this case each term has the common ratio \(\frac{1}{2}\) multiplied by it each time.